# DZ-TDPO: Non-Destructive Temporal Alignment for Mutable State Tracking in Long-Context Dialogue

### v1

@@ C.1 Selected Test Cases
Result: DZ-TDPO successfully navigates this by attending to the instruction "revert," showing that the learnable bias $\lambda$ does not destroy semantic understanding.
@@ C.2 Generation Metrics
Table 10: Generation Quality Metrics (Swapped View)
@@
Base Model 1.48 13.89 74.72
Standard DPO 0.66 11.2 74.25
SimPO 1.76 14.5 75.04
TDPO-DKL (w/o DZ-TA) 0.78 12.32 74.34
DZ-TDPO (Ours) 1.43 13.19 74.31
As presented in Table 10, we observe an intriguing divergence between n-gram metrics and alignment performance. SimPO achieves the highest scores across all generation metrics (e.g., ROUGE-L 14.50), yet suffers from a poor in-domain Win Rate (60.6%) and OOD generalization (30.8%). This phenomenon suggests "Superficial Mimicry": the model learns to imitate the surface-level lexical patterns of the reference responses without capturing the underlying state-tracking logic.
In contrast, DZ-TDPO achieves a slight decrease in n-gram overlap compared to the Base Model (ROUGE-L 13.19 vs. 13.89). We attribute this to the "Correctness-Mimicry Trade-off." Standard metrics like BLEU and ROUGE penalize any deviation from the reference. However, in temporal conflict scenarios, the model must deviate from historical patterns (which the Base Model tends to repeat) to assert the updated state. Our Adaptive Tau mechanism encourages such context-aware deviations, resulting in responses that are logically superior (SOTA Win Rate) despite having lower lexical overlap with static references. Meanwhile, Standard DPO shows a collapse in both quality metrics (BLEU 0.66) and stability, confirming the destructive nature of static constraints.
@@ C.3 Stress Testing under Massive Contextual Repetition
Accuracy 36% 78%
Inertia Failure 64% 21%
@@
The Base Model succumbs to the "frequency bias." Despite the instruction to find the final value, the sheer volume of historical tokens repeating the Old Value dominates the softmax attention calculation. The model effectively gets "brainwashed" by the repetition (64% failure rate).
DZ-TDPO achieves a 2.1x improvement in accuracy (78%). The DZ-TA mechanism ($\lambda\approx 0.68,\tau=20.0$) effectively penalizes the attention scores of the repeated historical tokens based on their distance.
@@ Appendix D Sensitivity Analysis
Small $\tau$ (e.g., $\tau=2$): The model becomes "myopic." While it achieves 87%+ win rates on immediate conflicts, it fails to maintain coherence across session boundaries, causing a slight rise in PPL.
Large $\tau$ (e.g., $\tau=32$): The mechanism degrades towards standard DPO. The "Historical Inertia" returns, and the win rate drops to  65%.
@@
2.0 87.9% 36.5 Aggressive Recency
4.0 87.3% 28.9 Balanced
8.0 (Ours) 86.2% 24.8 Optimal Trade-off
16.0 72.5% 24.2 Conservative
32.0 65.3% 23.9 Standard DPO
@@ Appendix E Computational Efficiency Analysis
Compared to standard DPO, DZ-TDPO increases training time by only  16% (due to the dynamic coefficient calculation and all-MiniLM-L6-v2 model). Peak VRAM usage remains identical as no new large matrices are introduced.
@@
Training Time 0.52 hours 0.6 hours +15.2%
Peak VRAM 56.5 GB 69.1 GB +22.3%
@@
Appendix F Visualization of Learned Attention Bias
@@
Initial State: The bias imposes a standard exponential decay.
Learned State: After training on MSC, the parameter $\lambda$ converged to approximately $0.68$ (from $0.5$). This indicates that the model actively *increased* the penalty on distant history to resolve conflicts, validating our hypothesis that standard models "under-decay" historical information.
Figure 2: Visualization of the Learned DZ-TA Mechanism. The plot compares the attention logit bias at initialization ($\lambda=0.5$, gray dashed) versus the learned state after training on MSC ($\lambda\approx 0.68$, red solid). The steeper slope in the learned curve indicates that the model actively optimized the parameter to impose a stronger penalty on distant history, thereby effectively suppressing outdated information to resolve temporal conflicts.

### v2

@@ C.1 Selected Test Cases
Result: DZ-TDPO successfully navigates this by attending to the instruction "revert," showing that the structural bias $\lambda$ does not destroy semantic understanding.
@@ C.2 Generation Metrics
Table 11: Generation Quality Metrics
@@
Base Model 1.23 13.22 74.95
Standard DPO 1.21 13.11 74.93
SimPO 1.07 13.08 74.80
TDPO-DKL (w/o DZ-TA) 1.02 12.92 74.34
DZ-TDPO (Ours) 1.10 13.03 74.40
As presented in Table 11, we observe an intriguing divergence between n-gram metrics and alignment performance. The Base Model achieves the highest scores across generation metrics (e.g., ROUGE-L 13.22), reflecting its strong tendency to mimic historical patterns and reference texts.
In contrast, DZ-TDPO exhibits a slight decrease in n-gram overlap compared to the Base Model (ROUGE-L 13.03 vs. 13.22). We attribute this to the "Correctness-Mimicry Trade-off." Standard metrics like BLEU and ROUGE penalize any deviation from the reference. However, in temporal conflict scenarios, the model must actively deviate from outdated historical contexts (which the Base Model tends to repeat) to assert the updated state. Thus, the marginal drop in lexical overlap, when combined with our superior Win Rate (55.4%), confirms that DZ-TDPO prioritizes logical correctness over surface-level mimicry. Meanwhile, other alignment baselines like SimPO and Standard DPO show similar trends, further validating that effective state tracking requires breaking away from the pre-trained manifold’s inertia.
@@ C.3 Stress Testing under Massive Contextual Repetition
Accuracy 27% 78%
Inertia Failure 73% 22%
@@
The Base Model succumbs to the "frequency bias." Despite the instruction to find the final value, the sheer volume of historical tokens repeating the Old Value dominates the softmax attention calculation. The model effectively gets "brainwashed" by the repetition (73% failure rate).
DZ-TDPO achieves a 2.89x improvement in accuracy (78%). The DZ-TA mechanism ($\lambda\approx 0.5,\tau=8.0$) effectively penalizes the attention scores of the repeated historical tokens based on their distance.
@@
C.4 Head-to-Head Comparative Evaluation
Beyond the win rates reported in Table 1 (calculated against the reference answers), we conducted direct pairwise battles between DZ-TDPO and the baselines on the MSC test set (N=500). As shown in Table 13, DZ-TDPO consistently outperforms both Standard DPO and SimPO, demonstrating that our method generates responses that are qualitatively preferred by the judge in direct comparisons.
Comparison Win Tie Loss Win Rate
vs. Standard DPO 289 8 203 58.6%
vs. SimPO 278 10 212 56.6%
Table 13: Head-to-Head evaluation results.
C.5 Visualization of Alignment Stability
To intuitively demonstrate the “Alignment Tax,” we visualize the training dynamics of validation perplexity (PPL) in Figure 2. Based on the starting checkpoint ($PPL\approx 22.1$) and the final converged states reported in Table 1, we reconstructed the divergence trends.
Figure 2: Illustration of Stability Dynamics. The plot visualizes the divergence in perplexity. Standard DPO (Red) exhibits a characteristic distribution collapse, where the rigid historical constraints force the model to degrade its linguistic capabilities to satisfy the reward ($PPL\rightarrow 102.3$). In contrast, DZ-TDPO (Blue) maintains a stable trajectory ($PPL\rightarrow 26.0$), confirming that our structural bias effectively decouples conflict resolution from general linguistic modeling.
This visual comparison reinforces our core claim: purely optimization-based methods (Standard DPO) struggle to reconcile state updates with historical priors, leading to catastrophic forgetting of the language manifold. DZ-TDPO’s “Dual-Zone” design acts as a stabilizer, allowing the model to learn the specific state update without sacrificing global coherence.
@@ Appendix D Sensitivity Analysis
Small $\tau$ (e.g., $\tau=2$): The model becomes "myopic." While it achieves 56%+ win rates on immediate conflicts, it fails to maintain coherence across session boundaries, causing a slight rise in PPL.
Large $\tau$ (e.g., $\tau=32$): The mechanism degrades towards standard DPO. The "Historical Inertia" returns, and the win rate drops to  47%.
@@
2.0 56.0% 39.5 Aggressive Recency
4.0 55.6% 32.9 Balanced
8.0 (Ours) 55.4% 26.0 Optimal Trade-off
16.0 50.2% 25.2 Conservative
32.0 47.6% 24.8 Standard DPO
@@ Appendix E Computational Efficiency Analysis
Compared to standard DPO, DZ-TDPO increases training time by only  15.4% (due to the dynamic coefficient calculation and all-MiniLM-L6-v2 model). Peak VRAM usage remains identical as no new large matrices are introduced.
@@
Training Time 0.52 hours 0.6 hours +15.4%
Peak VRAM 68.9 GB 69.1 GB +0.3%
@@
Appendix F Visualization of Structural Attention Bias
@@
Our preliminary sensitivity analysis indicated a potential optimal range around $\lambda\in[0.6,0.7]$. However, to prioritize regularization and prevent ’lazy head’ overfitting, we conservatively fixed $\lambda=0.5$. As shown in the visualization, even this baseline decay strength effectively suppresses outdated information.
Figure 3: Visualization of the Structural Attention Bias. The plot compares the attention logit bias of the standard decay ($\lambda=0.5$, gray dashed) versus a hypothetical aggressive decay ($\lambda=0.68$, red solid). The steeper slope of the hypothetical curve illustrates the stronger penalty required if we were to prioritize conflict resolution over stability; however, our fixed prior ($\lambda=0.5$) strikes a balanced trade-off, effectively suppressing outdated information on distant history to resolve temporal conflicts.
Appendix G LLM-as-a-Judge Evaluation Details
To ensure reproducibility, we provide the detailed configuration for the automated evaluation using DeepSeek v3.2.
To address position bias, our evaluation script employs a randomized shuffling mechanism. For every sample pair $(y_{\text{ours}},y_{\text{base}})$, we flip a fair coin ($p=0.5$) to determine which model is presented as Assistant A or Assistant B. The final verdict is mapped back to the original model identities.
The evaluation uses a strict instruction-following prompt. As shown in Table 16, the prompt is designed to enforce impartiality and adherence to specific evaluation criteria.
While DeepSeek-V3.2 demonstrates high correlation with human judgment, automated judges may exhibit verbose bias (preferring longer answers) or safety bias (preferring refusals). Although we implemented position swapping to mitigate bias, subtle semantic nuances in State Updates might still be misinterpreted by the judge. Future work involves verifying these results with human annotation on a subset of the data.
Component Content / Template
System Prompt You are a helpful assistant that acts as an impartial judge to evaluate the quality of AI responses.
Task
Description
I want you to act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user question displayed below. You should choose the assistant that follows the user’s instructions and answers the user’s question better.
Evaluation Criteria Your evaluation should consider factors such as usefulness, relevance, accuracy, depth, creativity, and level of detail.
Input Slots [User Question]
{question}
[Assistant A]
{answer_a}
[Assistant B]
{answer_b}
Output
Requirement
Output your decision in a strict JSON format with two keys: “reason” and “winner”. The “winner” must be one of “A”, “B”, or “Tie”.
Table 16: Evaluation Prompt Template. The structured prompt used for the automated judge. The input slots are dynamically populated with the randomized model outputs.
G.1 Pilot Human Validation
To validate the reliability of our automated judge (DeepSeek-V3.2), we conducted a meta-evaluation on a random subset of 51 samples. An expert annotator manually reviewed the preference pairs to establish a gold standard.
The analysis reveals a high degree of alignment between the automated judge and the human expert. Specifically, DeepSeek-V3.2 achieved an accuracy of 88.2% and a Cohen’s Kappa coefficient of 0.78, indicating "substantial agreement" (Landis and Koch, 1977). This confirms that the automated judge accurately captures the nuances of temporal state conflicts and serves as a reliable proxy for human preference (Zheng et al., 2024) in our experiments.
Furthermore, the confusion matrix (Table 17) reveals that the automated judge exhibits a conservative tendency. In 3 cases where the human expert judged DZ-TDPO as the winner (Win), the model judged it as a loss (Loss). Conversely, the model rarely "hallucinates" a win (only 1 case where Human=Loss but Model=Win). This suggests that our SOTA results reported in the main paper are likely lower-bound estimates, as the automated judge is stringent in awarding wins.
Table 17: Confusion Matrix of Human-Model Agreement. Rows represent the human expert labels; columns represent the automated judge (DeepSeek-V3.2). Labels: 0 (Loss), 1 (Tie), 2 (Win).
Automated Judge
Human Expert Loss (0) Tie (1) Win (2)
Loss (0) 17 0 1
Tie (1) 1 2 1
Win (2) 3 0 26